Using these data from the comparative balance sheet of Rollaird Company, perform vertical analysis. (Round percentages to 1 decimal place, e.g. 12.5%.)

Dec. 31, 2022

Dec. 31, 2021

Amount

Percentage

Amount

Percentage

Accounts receivable

$ 460,000

Enter a value in percentages rounded to 1 decimal place.

% $ 400,000

Enter a value in percentages rounded to 1 decimal place.

%

Inventory

$ 780,000

Enter a value in percentages rounded to 1 decimal place.

% $ 650,000

Enter a value in percentages rounded to 1 decimal place.

%

Total assets

$3,164,000

Enter a value in percentages rounded to 1 decimal place.

% $2,800,000

Enter a value in percentages rounded to 1 decimal place.

%
Answer #1
Dec. 31, 2022 Dec. 31, 2021
Amount Percentage Amount Percentage
Accounts receivable 460000 14.5% 400000 14.3%
Inventory 780000 24.7% 650000 23.2%
Total assets 3164000 100.0% 2800000 100.0%
Workings:
Dec. 31, 2022:
Accounts receivable 14.5% =460000/3164000
Inventory 24.7% =780000/3164000
Dec. 31, 2021:
Accounts receivable 14.3% =400000/2800000
Inventory 23.2% =650000/2800000